OpenStudy (anonymous):

The sum of the first 17 terms of an arithmetic sequence is 1,360. If the first term is 16, find the 17th term. Please help!

OpenStudy (anonymous):

Here is the equation you will use. S = sum n = number of terms (17), and a = first term (16) and b = 17th term\[\S = 1/2n(a+b)\]

OpenStudy (anonymous):

well the squiggly at the front is the S

OpenStudy (anonymous):

sum of 17 terms =17/2 [32 + 16d] = 1360 32 + 16d = 1360 * 2/17 d = 8 17th term = 16 + 16d = 144

OpenStudy (anonymous):

http://mathguide.com/lessons/SequenceArithmetic.html#calculate here is a website that describes this in detail. I see what jimmy is doing, but I am not entirely sure it applies to your problem situation.

OpenStudy (anonymous):

or atleast not by what is on this site. I remember using the equation I found.

OpenStudy (anonymous):

Hmmm... O.k. I'll check it out. Thank you so much guys! :)

OpenStudy (anonymous):

in there is a quicker way which i've just remembered sum = (n/2) [ a + l] where l = last term - in this case l = 17th term